OpenStudy (anonymous):

Rewrite the expression ln 9 + 3ln x + 2ln (x^2+4) as a single logarithm \ln A. Then the function A=

Nnesha (nnesha):

familiar with the `ln` rules?

OpenStudy (solomonzelman):

And there they are, the rules... ■ \(\color{#000000 }{ \displaystyle a\ln(b)=\ln(b^a) }\) ■ \(\color{#000000 }{ \displaystyle \ln(c)+\ln(d)=\ln(c\cdot d) }\)

Nnesha (nnesha):

thanks i forgot em. i need to focus on sclwork